OpenStudy (anonymous):

A 25-inch television has a screen that is 15 inches high. How wide is the screen?

OpenStudy (ddcamp):

Television screen sizes are the lengths of the diagonals. If the screen is 15 inches high and it's a 25-inch telivision, we have something like this: 

OpenStudy (anonymous):

I honestly forgot what to do next lol?

OpenStudy (ddcamp):

You know the Pythagorean theorem?

OpenStudy (anonymous):

a^2 + b^2=c^2?

OpenStudy (ddcamp):

Yup. Since the tv is a rectangle, the diagonal forms a right triangle. If c=25 and a=15, can you solve for b? 15²+b²=25²

OpenStudy (anonymous):

okay, I am going to solve it real quick. Let me know if I am wrong.

OpenStudy (anonymous):

I got 20

OpenStudy (ddcamp):

Yup, that's right

OpenStudy (trantom):

It should be 20, because yea.

OpenStudy (trantom):

xD

OpenStudy (anonymous):

lol @ trantom

OpenStudy (anonymous):

Idk how I forgot how to do this...

OpenStudy (ddcamp):

One thing you'll notice is that there are common patterns in right triangles. This is a 3-4-5 right triangle, because the ratios of the sides are 3:4:5

OpenStudy (anonymous):

Thank you so much though!

OpenStudy (trantom):

The pythagorean therom was one of the first things I learned in algebra. I learned it before probablility :P

OpenStudy (trantom):

*probability

OpenStudy (anonymous):

same, I just suck at remembering that solution..